OpenStudy (specek18):

What are the possible values of x in x2 + 3x + 3 = 0?

OpenStudy (godlovesme):

is that x^2?

OpenStudy (specek18):

yes. sorry posting a picture of the question!

OpenStudy (godlovesme):

thats alright but i think its b or c sorry im bad at radicals and stuff like this =0

OpenStudy (specek18):

it's okay..do you know anyone on here that could help?

OpenStudy (godlovesme):

@ageta can u please help :)

OpenStudy (specek18):

thank you for your help @godlovesme

OpenStudy (godlovesme):

np :)

OpenStudy (ageta):

sorry idk

OpenStudy (specek18):

it's okay. thank you. just really need help

OpenStudy (godlovesme):

do u know someone that could help @ageta

OpenStudy (specek18):

@TheSmartOne @sleepyjess @e.mccormick @Sheraz12345 @magepker728

OpenStudy (ageta):

Let's solve your equation step-by-step. x2+3x+3=0 Step 1: Use quadratic formula with a=1, b=3, c=3. x=−b±sqrt(b2−4ac) 2a x=−(3)±sqrt((3)2−4(1)(3))/2(1) x=−3±sqrt(−3)/2

OpenStudy (sheraz12345):

Here he did it ^ Lol I was gonna say the same anyways Lol

OpenStudy (ageta):

does that help

OpenStudy (specek18):

@ageta no. i can't see it:( whats the answer?

OpenStudy (sheraz12345):

Its b

OpenStudy (ageta):

x=−3±sqrt(−3)/2

OpenStudy (specek18):

@Sheraz12345 for real?

OpenStudy (specek18):

@ageta do you agree it's the second? sorry just gotta get it right

OpenStudy (sheraz12345):

no sorry

OpenStudy (ageta):

no it is C

OpenStudy (sheraz12345):

Yea its C

OpenStudy (specek18):

okay...so for sure it's c?

OpenStudy (ageta):

yes

OpenStudy (specek18):

okay thank you.

OpenStudy (12man):

it is A

OpenStudy (ageta):

no it is C

OpenStudy (12man):

i meant C

OpenStudy (specek18):

here's my last question. will give medals when i find out if they're right after this

OpenStudy (ageta):

x=±sqrt(−54) is x^+58=4 ..

OpenStudy (specek18):

i can't see your answers...

OpenStudy (ageta):

x=±sqrt(−44) is x^2+47=3

OpenStudy (ageta):

x=±sqrt(−23) is 2x^2+50=4

OpenStudy (specek18):

ugh. it's showing up as all question marks. try taking a screen shot before you press post and just try to post it as the screenshot/ picture

OpenStudy (ageta):

sorry

OpenStudy (ageta):

x=±sqrt(−39) is 2x^2+85=7

OpenStudy (specek18):

i can't see it;/ i'm sorry. can you email it to me? limegirl4me@aol.com

OpenStudy (12man):

x=sqrt(39) is 2x^2+85=7

OpenStudy (12man):

x=sqrt(9) is 2x^2+85=7

OpenStudy (12man):

x=sqrt(39) is 2x^2+85=7

OpenStudy (12man):

x=sqrt(23) is 2x^2+50=4

OpenStudy (specek18):

thank you...but you used one option for two...

OpenStudy (ageta):

wait

OpenStudy (specek18):

i didn't enter it yet...

OpenStudy (ageta):

x=±sqrt(−39)2is x^2+85=7

OpenStudy (specek18):

@ageta message it to me. i can't see it

OpenStudy (12man):

x=sqrt(39)2is x^2+85=7

OpenStudy (ageta):

x=sqrt(23) is 2x^2+50=4

OpenStudy (specek18):

two left:) and thank you so much! i can see them. just type the last two one at a time please:)

OpenStudy (ageta):

x=±sqrt(11) is 2x2+50=4

OpenStudy (specek18):

@ageta do u know the last two?

OpenStudy (ageta):

and the last one is the others

OpenStudy (specek18):

i failed. but thanks anyways.